### v1

# DeepExtractor: Time-domain reconstruction of signals and glitches in gravitational wave data with deep learning

Abstract
Gravitational wave (GW) interferometers, such as LIGO, Virgo, and KAGRA, detect faint signals from distant astrophysical events, such as binary black hole mergers. However, their high sensitivity also makes them susceptible to background noise, which can obscure these signals. This noise often includes transient artifacts called “glitches”, that can mimic genuine astrophysical signals or mask their true characteristics. Fast and accurate reconstruction of both signals and glitches is crucial for reliable scientific inference. Accurate GW signal reconstruction facilitates parameter estimation of the GW source, while glitch reconstruction—challenging due to the unmodeled nature of glitches—supports glitch mitigation and enables mock data challenges involving realistic noise artifacts. In this study, we present DeepExtractor, a deep learning framework that is designed to reconstruct signals and glitches with power exceeding interferometer noise, regardless of their source. We design DeepExtractor to model the inherent noise distribution of GW interferometers, following conventional assumptions that the noise is Gaussian and stationary over short time scales. It operates by predicting and subsequently subtracting the noise component of the input mixture of noise and signal or glitch, retaining only the clean reconstruction. Our innovative approach achieves superior generalization capabilities for arbitrary signals and glitches compared to methods that directly map mixture inputs to the clean training waveforms, yielding precise reconstruction of time-domain signals and glitches. We focus on applications related to glitches and validate DeepExtractor’s effectiveness through three experiments: (1) reconstructing simulated glitches injected into detector noise, (2) comparing its performance with the state-of-the-art BayesWave algorithm, and (3) analyzing real data from the Gravity Spy dataset to demonstrate effective glitch subtraction from LIGO strain data. Our proposed model achieves a median mismatch of only $0.9\%$ for simulated glitches, outperforming several deep learning baselines. Additionally, DeepExtractor surpasses BayesWave in glitch recovery, offering a dramatic computational speedup by reconstructing one glitch sample in approximately $0.1$ seconds on a CPU, compared to BayesWave’s processing time of approximately one hour per glitch.
preprint: APS/123-QED
I Introduction
The first detection of gravitational waves (GWs) from a binary black hole (BBH) merger in 2015 ushered in a new era of astrophysics [1]. Since then, Advanced LIGO [2] and Advanced Virgo [3] detectors have made confident detections of $\sim 90$ compact binary coalescence (CBC) GW events [4, 5, 6], accumulated over three observing runs (O1, O2, O3). With the fourth observing (O4) run currently underway, hundreds of more detections are expected from the enhanced sensitivity of GW detectors [7] and the introduction of KAGRA [8], Japan’s underground detector.
GW data analysis procedures, such as those regarding the estimation of GW source parameters [9, 10, 11], typically assume the detector noise to be stationary and Gaussian [12]. This assumption breaks down in the presence of ‘glitches’; transient noise artifacts that introduce non-Gaussian and non-stationary features in the data [13, 14, 15]. Glitches are unmodeled terrestrial noise events stemming from environmental factors (e.g., earthquakes, wind, anthropogenic noise) or instrumental issues (e.g., control systems, electronic components [16]). However, the origins of many glitches remain unknown [12].
Model-free algorithms, Omicron [17] and Gravity Spy [18], are utilized for glitch detection and characterization, respectively. These tools visualize and identify glitches by analyzing excess power in Q-scans (spectrograms) [19], which are time-frequency representations. Glitches exhibit a broad range of time-frequency morphologies, which Gravity Spy classifies into over 20 distinct categories. Figure 1 illustrates a representative glitch from the ‘tomte’ class, showing its time-series data alongside the corresponding Q-scan used in Gravity Spy analysis.
Glitches can resemble astrophysical signals, increasing false positives  [20, 21, 22, 23, 24, 25, 26, 27, 28] and hindering subsequent analyses. However, when glitches overlap with a GW signal they can significantly bias parameter estimation [29, 30, 31] and must be carefully modeled and removed before the GW signal can be used to make any measurements. This is a commonplace occurrence, with $\sim 25$ out of the $\sim 90$ confident detections so far requiring some form of glitch mitigation [6, 29, 32, 33, 34, 35].
Ongoing upgrades to current detector systems and the introduction of third-generation detectors, such as the Einstein Telescope (ET) [36] and Cosmic Explorer (CE) [37], will significantly increase the detection rates from all sources of GWs. For instance, it is estimated that the ET alone will detect on the order of $\mathcal{O}(10^{4})\,\text{y}^{-1}$ BBH mergers and binary neutron star (BNS) mergers [38, 39]. This could lead to over 400 GW events per day. If glitch rates remain similar (approximately 1 per minute during the third observing run [6, 5]), higher detection rates will naturally exacerbate issues relating to glitches [7]. Such challenges demand fast reconstruction algorithms that can accurately separate events from the underlying detector background, agnostic of their sources.
(a) Time series
(b) Q-scan
Figure 1: An example of a ‘tomte’ glitch in LIGO Livingston at GPS time $1263921313$ with an signal-to-noise ratio (SNR) of $48.6$.
In this study, we present DeepExtractor, a power spectral density (PSD)-informed deep learning framework designed to accurately reconstruct power excesses above a Gaussian detector background in the time domain. By simultaneously analyzing magnitude and phase spectrograms from the short-time Fourier transform (STFT), DeepExtractor isolates the underlying noise from signal and glitch events using a U-Net architecture [40]. The signal or glitch is then reconstructed by subtracting the predicted noise, following an additive model. Assuming additive interactions and linear approximations of signal and glitch components using time-domain wavelets, we simulate five waveform classes. By learning to isolate background noise from these classes, DeepExtractor is capable of interpolating to arbitrary signals and glitches.
DeepExtractor demonstrates superior accuracy in reconstructing simulated glitches and generalizes well to unseen distributions, outperforming several deep learning models. We also present a comparison of glitch reconstructions by DeepExtractor to those obtained using the state-of-the-art BayesWave [41], an advanced algorithm used in GW data analysis to model and subtract glitches from the strain data. Finally, we showcase DeepExtractor’s effectiveness in reconstructing real Gravity Spy [18] glitch events from LIGO’s third observing run, verified through Q-scans and time series plots of the detector data before and after glitch subtraction. This marks the first comprehensive application of deep learning for such tasks in GW data analysis.
This paper is structured as follows; in section II, we discuss the state-of-the-art in glitch modelling and mitigation in LIGO and Virgo and state-of-the-art machine learning research in time series denoising. In section III, we discuss our method, which includes an overview of our framework, our model architecture, the generation of training data, and the experiments implemented to show our DeepExtractor’s efficacy. In section IV, we present the results of the experiments and section V discusses the conclusions of our research.
II Related Work
II.1 GW glitch mitigation
A variety of methods have been developed to model and subtract glitches from the strain data that leverage machine learning [42] and parametric [43, 44] models. However, up to O3, BayesWave [41, 45, 29, 46] and gwsubtract [47, 25] [48] have been the only algorithms used for modeling and subtracting glitches in LIGO-Virgo-KAGRA data analyses [48]. While gwsubtract relies on strongly correlated auxiliary witness sensors to linearly subtract glitches from the strain data, BayesWave models glitches using the detector strain only.
BayesWave models the detector strain $h(t)$ in each interferometer as a linear combination:
$h(t)=n(t)+s(t)+g(t),$ (1)
where $n(t)$ represents Gaussian noise, $s(t)$ an astrophysical signal, and $g(t)$ a glitch. Non-Gaussian components, including signals and glitches, are represented as sums of sine-Gaussian (Morlet-Gabor) wavelets:
$\Psi(t;\vec{\theta})=Ae^{-(t-t_{0})^{2}/\tau^{2}}\cos\left(2\pi f_{0}(t-t_{0})+\phi_{0}\right),$ (2)
where $\vec{\theta}=\{t_{0},f_{0},Q,A,\phi_{0}\}$ denotes the wavelet parameters: central time ($t_{0}$), central frequency ($f_{0}$), quality factor ($Q$), amplitude ($A$), and phase offset ($\phi_{0}$) with $\tau=Q/2\pi f_{0}$. BayesWave performs Bayesian inference to estimate the wavelet parameters and the total number of wavelets of a glitch using a Reversible-Jump Markov Chain Monte Carlo (RJMCMC) sampler, which allows for the flexibility of models with variable dimensions.
For the glitch model, wavelet parameters are independent across detectors. For the signal model, a single set of wavelets is projected onto detectors based on extrinsic parameters such as the source sky location and polarization. The PSD is modeled with a combination of a cubic spline and Lorentzians, using the method from [49].
Three model assumptions guide the reconstruction of transient features in the data:
  1. 1.
    Gaussian noise with an astrophysical signal.
  2. 2.
    Gaussian noise with an instrumental glitch.
  3. 3.
    Gaussian noise with both a signal and a glitch.
Glitch subtraction uses either the glitch-only model or the signal+glitch model. The glitch-only model applies when the glitch is temporally or spectrally distinct from the signal. The signal+glitch model, used for significant overlaps, relies on data from multiple detectors to separate coherent signal power from incoherent glitch power, ensuring the astrophysical signal is preserved.
BayesWave outputs a posterior distribution of the model parameters, derived from the RJMCMC. A random sample from this posterior distribution is used to reconstruct the time domain glitch, which is subtracted from the strain data to mitigate the glitch [46].
II.2 Deep learning for signal denoising
Reconstructing GWs and glitches parallels time-series denoising tasks in fields like speech enhancement, acoustic processing, and audio source separation, all of which have seen significant advancements. Traditional methods such as dictionary learning, wavelet shrinkage, and singular value decomposition (SVD) have been effective [50, 51, 52]. However, supervised deep learning has revolutionized performance, particularly in data-rich applications like image denoising and speech enhancement [53, 54].
Time vs. time-frequency domains: Networks for denoising can process data in the time or time-frequency domain, each with trade-offs [55]. Time-domain approaches analyze raw waveforms directly, leveraging deep learning’s ability to extract features automatically [56, 57, 58, 59, 60]. Conversely, time-frequency representations like the STFT offer detailed signal characterizations, enhancing learning [61, 62, 54, 63]. However, STFT-based methods require pre- and post-processing steps, increasing computational load [64]. Key STFT parameters, such as frame size and overlap, affect time and frequency resolution but are often fixed rather than optimized jointly with network parameters.
Phase considerations in STFT: Most STFT-based denoising models use only the magnitude spectrogram, assuming the mixture (input) and source (target) phases are equivalent [65]. While effective, this assumption is not always valid, especially in signals with overlapping harmonics and multiple studies have shown the importance of phase for denoising performance [66, 67]. Iterative methods like Griffin-Lim [68] can approximate the source phase but are computationally intensive and may yield suboptimal results [69]. Recent approaches explore complex spectrograms that preserve phase or estimate it directly, showing improved performance [70, 71, 72].
Deep learning architectures for denoising: Popular architectures for denoising include convolutional neural networks (CNNs), convolutional autoencoders, and U-Nets. CNNs have been applied to tasks such as electrocardiogram denoising, speech enhancement, image denoising and source separation in time-frequency domains [73, 74, 53, 75]. Convolutional autoencoders compress signals into a latent representation and reconstruct them, achieving success in areas such as speech enhancement, health management and fault diagnosis [76, 77, 78, 79, 80]. However, deep autoencoders risk information loss in the compressed layer, especially with increasing the number of layers [81].
U-Nets [40] mitigate this issue with skip connections that retain features across layers, achieving superior results in tasks like acoustic source separation and image denoising [82, 58, 83]. Their ability to combine feature preservation with a large receptive field has made them a state-of-the-art choice for signal denoising [84].
Figure 2: An overview of DeepExtractor’s reconstruction approach. Two seconds of detector strain $h(t)$ is processed into magnitude and phase spectrogramns using an STFT. This is fed through our network, which maps the instance to the magnitude and phase spectrograms of the underlying background noise. We then use an inverse STFT to yield to corresponding time series for $n(t)$. An estimation of the underlying signal or glitch can be simply calculated as $h(t)-n(t)$.
II.3 Deep denoisers for GW data analysis
Recent studies have explored deep learning for denoising GW signals, focusing on isolating astrophysical events from noise in LIGO and Virgo data with promising results across various scenarios.
Recurrent autoencoders based on Long-Short-Term Memory (LSTM) networks [85, 86] have demonstrated robust performance in recovering GW signals. Trained on simulated Gaussian noise, these networks excel in real, non-Gaussian LIGO noise, outperforming traditional methods like principal component analysis and dictionary learning, especially at low signal-to-noise ratio (SNR). Notably, the model generalizes well to eccentric GWs outside the training set, showcasing adaptability in diverse signal environments. Autoencoders using CNNs in the encoder and LSTMs in the decoder have been applied in the AWaRe framework to reconstruct GW events that align with benchmark algorithms. The framework also incorporates uncertainty estimation and can reconstruct signals that overlap with glitches [87, 88].
WaveNet architectures [89] have also been employed for GW denoising in simulated Gaussian and real non-Gaussian, non-stationary LIGO noise, achieving high overlaps ($O>0.97$) between predicted and numerical relativity waveforms. These networks can additionally extract sine-Gaussian glitches due to structural similarities with GWs while remaining resilient to dissimilar glitches like Gaussian pulses.
U-Net models have been applied to denoise BBH signals using time-frequency representations [90]. Adapted from seismic signal separation studies [91], this approach leverages the real and imaginary parts of spectrograms to predict separate masks for signal and noise. These masks, applied to input spectrograms, allow reconstruction of the time-domain signal via inverse STFT, yielding strong performance on real astrophysical events.
NNETFIX [92] addresses signal reconstruction when portions are gated due to glitches. Simulating glitches of varying durations and offsets, this method effectively reconstructs BBH signals with single-interferometer SNRs above $20$, even for glitches lasting hundreds of milliseconds near merger times. This highlights its relevance in the upcoming LIGO-Virgo-KAGRA observing run, where glitch overlap with longer-duration events like BNS signals is expected to increase.
III Methods
III.1 DeepExtractor
III.1.1 Overview of framework
DeepExtractor is a deep learning framework designed to reconstruct signals or glitches in GW data with power exceeding the underlying detector noise. An overview of the framework is illustrated in Figure 2. Built using PyTorch’s deep learning library [93], DeepExtractor’s codebase is publicly accessible on GitLab [^1] https://git.ligo.org/tom.dooney/deepextractor.git. DeepExtractor is agnostic to the origin and morphology of excess power in the detector, whether it is a signal or a glitch. Throughout this paper, we use $g(t)$ to denote either a signal or a glitch for convenience, given our primary focus on glitch-related applications. Although we often refer to events as glitches, this is solely for convenience, as the model does not distinguish between the origins of excess power.
The framework operates under the first two models of BayesWave described in Section II.1, where the data contains either an astrophysical signal or a glitch. These models correspond to BayesWave’s signal-only or glitch-only modes. Currently, DeepExtractor does not handle the third model assumption, where a glitch overlaps with an astrophysical signal (BayesWave’s signal+glitch mode). While this capability is critical for a fully-functional glitch mitigation algorithm, this study focuses on the simpler signal-only and glitch-only cases as a foundational step. Potential approaches for addressing the signal+glitch scenario are discussed in Section V.
We follow BayesWave’s assumption in Eq. 1 that a glitch $g(t)$ interacts additively with the strain data in a real detector setting. Our goal is to map the input mixture $h(t)$ to the noise component $n(t)$, where $h(t)$ is represented as:
$h(t)=n(t)+g(t)$ (3)
We can then subtract the model output $\hat{n}(t)$ from $h(t)$ to yield an estimate of the underlying signal or glitch embedded in the detector background via $\hat{g}(t)=h(t)-\hat{n}(t)$.
III.1.2 Training strategy
The DeepExtractor framework utilizes a neural network that directly outputs the background noise component from the data in the spectrogram domain. Our core idea is that the background noise is easier to model by our network because it is simpler and more stationary than the diverse glitches.
[…]
DeepExtractor is trained to learn the mapping in the time-frequency domain using invertible STFT transforms. Modeling both magnitude and phase is essential for effective signal and glitch reconstruction in the time domain. To provide the inputs and targets, we first transform the time-domain signals $h(t)$ and $n(t)$ into the time-frequency domain by computing their complex spectrograms, $h(t,f)$ and $n(t,f)$, using the STFT (see Section III.2.1). These spectrograms are then decomposed into magnitude and phase components, which are concatenated channel-wise before being fed into the network. Thus, DeepExtractor takes the magnitude and phase components of $h(t,f)$ as input and outputs the corresponding components of the estimated noise $\hat{n}(t,f)$.
We optimize the network by minimizing a mean squared error (MSE) loss function, which compares the predicted components $\hat{n}(t,f)$ with the true background noise components $n(t,f)$, as defined below:
$\frac{1}{N}\sum_{i=1}^{N}\left(\hat{n}_{i}(t,f)-n_{i}(t,f)\right)^{2},$ (4)
where $i$ is the index of the data sample and $N$ is the batch size.
During inference, the predicted magnitude and phase components of $\hat{n}(t,f)$ are recombined into a complex spectrogram, which is then transformed back to the time domain using the inverse STFT to produce the noise estimate $\hat{n}(t)$. By subtracting $\hat{n}(t)$ from the input mixture $h(t)$, the framework isolates the underlying signal or glitch, $\hat{g}(t)$. This approach ensures the network focuses on learning the simpler distribution of Gaussian noise, avoiding the complexities associated with directly modeling the diverse combinations of glitches present in the data.
DeepExtractor employs the Adam optimizer [94] with a batch size of 32. Losses for DeepExtractor and other time-frequency models are calculated directly on spectrogram data, while for time-domain benchmarks, they are computed on raw time-series data. To optimize the learning process, the training incorporates the ReduceLROnPlateau scheduler provided by Pytorch, which reduces the learning rate after four epochs without improvement, and early stopping is applied after ten such epochs. This ensures an efficient yet flexible training duration tailored to the model’s performance on validation data.
III.1.3 U-Net Architecture
Figure 3: The DeepExtractor architecture implemented on STFT data. The two input and output channels process the magnitude and phase components of the STFT respectively.
Motivated by the discussion in Section II, DeepExtractor adopts a U-Net architecture for its ability to capture both high- and low-level features through skip connections. Skip connections concatenate encoder features with their corresponding decoder features, as illustrated in Figure 3. These connections preserve fine-grained details while enabling a large receptive field, which is particularly advantageous for analyzing spectrogram representations of signals or glitches with broad time-frequency ranges. U-Nets effectively reconstruct inputs without significant information loss.
DeepExtractor is a 2D model designed to process magnitude and phase spectrograms. It accepts two input channels and outputs two spectrograms of the same dimensionality, as illustrated in Figure 3.
Encoder (Downsampling Path): The encoder extracts hierarchical features from the input through:
  • DoubleConvolution Blocks: Each block consists of two 2D convolutional layers (kernel size 3), each followed by batch normalization [95] and ReLU activation [96], to enhance feature extraction while preserving spatial resolution.
  • MaxPooling: A 2D max-pooling operation (kernel size 2) reduces spatial resolution by half after each DoubleConvolution block, enabling hierarchical feature representation.
Feature maps from the encoder are stored as skip connections for the decoder.
Decoder (Upsampling Path): The decoder reconstructs the original input resolution while integrating features from the encoder:
  • Transposed Convolutions: Upsampling is performed using transposed convolutional layers that double the spatial resolution.
  • Skip Connections: Features from the encoder at corresponding resolutions are concatenated with the upsampled features to preserve fine details.
  • DoubleConvolution Blocks: Combined features are processed through DoubleConvolution blocks for refinement.
Bottleneck: At the deepest level, the bottleneck processes the feature map using a DoubleConvolution block, capturing the most abstract representations of the input.
Output: The model concludes with a single 2D convolutional layer (kernel size 1) with a linear activation function, producing outputs with the desired dimensionality.
Time-Domain Variant: We also experiment using a 1D time-domain counterpart of DeepExtractor’s U-Net architecture (UNET1D). This replaces all 2D operations with 1D equivalents and processes a single input/output channel. Specifically, the 1D model maps the raw, whitened time series $h(t)$ to the noise component $n(t)$, omitting the STFT and iSTFT transformations used in the 2D variant, as shown in Figure 2.
III.2 Datasets
(a) Chirp
(b) Sine
(c) Sine-Gaussian
(d) Gausian Pulse
(e) Ringdown
Figure 4: Examples of each of the five training glitch classes; chirp, sine, sine-gaussian, gaussian pulse and ringdown. For each 222\,2s training sample, anywhere between 1 and 30 of these signals (selected randomly) are injected into the Gaussian background sample.
III.2.1 Simulated data
To properly validate our model, we first implement a fully simulated approach for both training and validation, simulating the background noise and glitches. We simulate white noise training targets $n(t)$ with a duration of 2 seconds at a sampling rate of $4,096\,$ Hz, yielding 8,092 data points. The simulated noise has a mean of 0 and variance of $f_{s}/2=2,048\,$Hz, where $f_{s}$ is the sampling rate. We generate 250,000 background samples for training and 25,000 for validation.
To create the corresponding input samples, we inject various analytical waveforms, represented by $g(t)$, into the background noise $n(t)$, as per Eq. 3. These waveforms are drawn from five proxy glitch classes: chirp, sine, sine-Gaussian, Gaussian pulse, and ringdown, as shown in Figure 4. Each waveform class has its own parameters that are randomly sampled during generation, such as frequency range (from 111\,1 Hz to the Nyquist frequency of $2,048\,$Hz), bandwidth, and duration (randomly chosen between 0.1250.1250.125\,0.125s and 222\,2s). For more information on these training classes, please see Appendix VII.1 and our code repository. The SNR is varied between 1 and 250, scaled according to
[…]
where $g(f)$ and $S_{n}(f)$ represent the Fourier transform of the injected waveform and the detector noise PSD, respectively [97]. The scaling ensures that each waveform achieves the desired SNR within the specified range.
To improve generalization and avoid overfitting, we inject linear combinations of glitches into each background sample, instead of using a single glitch $g(t)$, where:
$\mathbf{g}=\sum_{i=1}^{k}g_{i}(t)$ (6)
and $k\in[1,30]$. Thus, $\mathbf{g}$ will replace $g(t)$ in Eq. 3 when describing the training data moving forward.
This allows for between 1 and 30 glitches per sample, sampled uniformly across the five classes. This follows BayesWave’s method of modeling signals and glitches as sums of sine-Gaussian wavelets (Eq. 6), but increases the space of analytical models with additional classes to improve the network’s generalization ability.
Additionally, glitches can be scaled to low SNRs, so some training samples include injections with very little contribution to the data. This approach helps the model to indirectly learn the background PSD, enhancing its capability to distinguish and reconstruct subtle glitch features while maintaining robustness across diverse detector noise characteristics.
To prepare the data for training DeepExtractor, we scale (standardize) the inputs and outputs using StandardScaler from sklearn [98], which transforms the data so that it has a mean of 0 and a standard deviation of 1. This standardization method assumes a Gaussian distribution of the data and is appropriate given the Gaussian-like characteristics of the background noise. We fit the scaler to the model input only where glitches are present, only transforming the background noise target with the input-fitted scaler. This ensures that the input and output occupy the same standardized space. This preprocessed data is used to train the time-domain benchmarks for DeepExtractor.
Short time Fourier transform (STFT): As mentioned in Section III.1.2, DeepExtractor and the 2D benchmarks are trained on spectrograms calculated from the above preprocessed data using the STFT. Since the time-domain data is already standardized, no additional scaling is performed on the spectrograms. Figure 5 shows examples of the magnitude and phase components for both the input and target spectrograms. We choose STFT parameters to ensure compliance with the Constant Overlap-Add (COLA) condition [99] (please see Appendix VII.2 for more information). A Hann window is employed to minimize spectral leakage, further improving the quality of the spectrogram representation [100]. These chosen settings yield complex spectrograms with dimensions of $257\times 257$, which we separate into magnitude and phase spectrograms. To investigate the effect of varying time and frequency resolutions on signal and glitch reconstruction, we also experiment with alternate parameter configurations that produce spectrograms of dimensions $129\times 129$ and $65\times 129$ while ensuring that all configurations satisfy the COLA condition.
(a) Magnitude
(b) Phase
Figure 5: Magnitude and phase STFT spectrograms for noise+glitch input (top) and noise-only target (bottom) used to train DeepExtractor.
III.2.2 Real LIGO O3 backgrounds
(a) PSD of detector noise in LIGO Hanford calculated over 512 seconds and centered at GPS time $1256665874$.
(b) PSD of the Hanford noise after whitening (blue) and simulated white noise (green) described in Section III.2.1.
Figure 6: Real vs simulated detector noise.
GW detector noise is often modeled as Gaussian and stationary over short timescales, but it is generally non-Gaussian and exhibits a time-varying PSD. The sensitivity of GW detectors is constrained by broadband noise sources, such as seismic noise at low frequencies and quantum sensing noise at high frequencies [12]. An example PSD from LIGO Hanford is shown in Figure 6. The bottom plot compares the PSD of whitened data using Welch averaging [101] to that of simulated white noise described in Section III.2. While whitening reduces broadband noise, narrow-band spectral lines—caused by electrical and mechanical sources or resonances—persist [102]. In contrast, the simulated noise PSD shows a smooth, flat trend with no spectral lines, underscoring the complexities of real GW detector noise.
To accurately reconstruct signals and glitches in real GW data, it is essential to account for these unique noise features. A model trained exclusively on simulated data will inadvertently suppress spectral lines when applied to real data, leading to contaminated reconstructions of both signals and glitches. Therefore, incorporating real data into the training of DeepExtractor is crucial for adapting to the non-Gaussian, non-stationary nature of GW detector noise and ensuring accurate modeling of real events outside the training distribution.
To address this challenge, we apply transfer learning [103] to our DeepExtractor model, leveraging knowledge learned from simulated data to enhance performance or reduce training effort on real data. Transfer learning allows a model trained for one task to adapt to a related task, making it an effective strategy in this context.
We restrict the scope of this study to LIGO data during O3. We accumulate real backgrounds by identifying GPS times for all GW events and Omicron triggers. Background segments of 141414\,14s are extracted from GWOSC [104] data during O3a and O3b, ensuring these segments are at least 222\,2s away from any trigger. These segments are whitened using Welch’s method [101], with the first and last 222\,2s discarded to mitigate boundary artifacts. The resulting 101010\,10s segments are divided into 222\,2s samples. We perform a final validation step using Q-scans for each $2\,\mathrm{s}$ sample, discarding any sample with a maximum pixel (tile) energy — defined as the squared magnitude of the Q-transform coefficient for a time-frequency tile — exceeds 20 (see [105] for details on pixel energy). This ensures that no sample contains excessive power that could interfere with the training of DeepExtractor. After applying this criterion, we collect a total of $45,400$ background samples from LIGO’s O3 run, distributed as follows: $18,400$ from Hanford O3a, $1,300$ from Livingston O3a, $18,600$ from Hanford O3b, and $7,100$ from Livingston O3b. [^2] While Livingston has nearly twice as many Omicron triggers as Hanford, a significant portion of Livingston background samples were also excluded during the Q-scan validation due to persistent excess power, even outside the Omicron triggers. For injections, we follow the approach described in Section III.2.1, with the key distinction that the injections are now added to real, whitened O3 backgrounds.
III.2.3 Unseen Test Classes
To evaluate the generalization performance of DeepExtractor, we utilize two glitch generators: gengli and cDVGAN [106, 107]. These generative deep learning models produce realistic glitches, also enabling robust testing on unseen glitch classes. Trained on blip glitches extracted from LIGO O2 data using BayesWave, gengli generates reliable representations of this specific glitch type. A conditional GAN trained on preprocessed LIGO O3 data, cDVGAN produces blip [^3] https://gswiki.ischool.syr.edu/en/Glitches/Blip and tomte [^4] https://gswiki.ischool.syr.edu/en/Glitches/Tomte glitches along with BBH signals with component masses in the range of $[30,160]M_{\odot}$, a spin of zero and fixing $m_{1}>m_{2}$. As a conditional model, cDVGAN can create hybrid samples that blend features across all three training classes, introducing even greater diversity. By leveraging these generative models, we create a diverse test dataset that mimics the variety and complexity of real detector glitches. This dataset provides a robust benchmark for assessing DeepExtractor’s ability to generalize to unseen classes.
III.3 Experiments
III.3.1 Simulated Experiments
We evaluate DeepExtractor’s performance by injecting simulated glitch samples into simulated background noise. This controlled environment allows for a direct comparison between the extracted glitches and their ground-truth counterparts, helping identify the most effective data representations in terms of dimensionality and domain (time vs. time-frequency).
Simulated glitches are generated using both analytical waveforms and samples from the glitch generators discussed in Section III.2.3. While the analytical waveforms represent known classes from training, we include unseen classes using glitches simulated with gengli [106] and cDVGAN [107]. Samples are injected into simulated white noise at SNRs $\in[7.5,100]$ using Eq. 5. We limit the maximum SNR of the test glitches to 100 (reduced from 250 in our training set) to create a controlled environment for glitch reconstruction. Furthermore, over 90% of glitches observed during O3 had an SNR below 100.
The results are evaluated using the mismatch metric, defined as
$\mathcal{M}(g_{1},g_{2})=1-M(g_{1},g_{2})$ (7)
where
$M(g_{1},g_{2})=\max_{\phi_{c},t_{c}}\frac{\langle g_{1}|g_{2}\rangle}{\sqrt{\langle g_{1}|g_{1}\rangle\langle g_{2}|g_{2}\rangle}}$ (8)
with the inner product given by:
$\langle g_{1}|g_{2}\rangle=4\,\Re\int_{f_{\text{low}}}^{f_{\text{high}}}\frac{\tilde{g}_{1}(f)\tilde{g}_{2}^{*}(f)}{S_{n}(f)}\,\mathrm{d}f$ (9)
and the normalization:
$\langle g|g\rangle=4\,\int_{f_{\text{low}}}^{f_{\text{high}}}\frac{|\tilde{g}(f)|^{2}}{S_{n}(f)}\,\mathrm{d}f$ (10)
Here, $\tilde{g}_{1}(f)$ and $\tilde{g}_{2}(f)$ are the Fourier transforms of the waveforms (* represents complex conjugate), $S_{n}(f)$ is the PSD of the noise and $f_{\text{low}}$ and $f_{\text{high}}$ are the frequency bounds of the analysis. The match is maximized over relative shifts in time $t_{c}$ and phase $\phi_{c}$.
We evaluate the performance of DeepExtractor by comparing it against several deep learning benchmarks. These include a 1D denoising CNN (DnCNN1D) [53] and 1D and 2D autoencoders [108] that are structurally similar to the U-Net, but without skip connections. We also assess the performance of the time-domain variant of DeepExtractor’s U-Net architecture (UNET1D) to compare its effectiveness across domains. This comparison helps evaluate DeepExtractor’s performance relative to other denoising and feature extraction architectures. Furthermore, we examine the performance of both DeepExtractor and UNET1D when directly mapping to $g(t)$, investigating the effectiveness of our approach in first mapping to $n(t)$ and subtracting it to yield $\hat{g}(t)$. Finally, we apply DeepExtractor to background samples without injections, where the ideal outcome is for the output to match the input, and the inferred glitch component $\hat{g}(t)$ to be zero.
We explore additional target mappings for UNET1D by conducting several ablation studies. An ablation study in machine learning is a systematic experimentation technique used to understand the contribution of individual components (target mappings) of a model to its overall performance.
Dual-Channel Output: We evaluate a dual-channel output configuration where the network predicts $\hat{n}(t)$ in one channel and $\hat{\mathbf{g}}(t)$ in the other, using both $n(t)$ and $\mathbf{g}(t)$ as targets. An extra loss ensures the sum of outputs equals the mixture input ($\hat{\mathbf{g}}(t)+\hat{n}(t)==h(t)$), enforcing consistency with Eq. 3.
Difference Output Layer: In a variation of the dual-channel setup, we omit the explicit $\mathbf{g}(t)$ target. Instead, the second channel predicts $\hat{\mathbf{g}}(t)$ indirectly as $\hat{\mathbf{g}}(t)=h(t)-\hat{n}(t)$. This approach has been successfully applied in speech enhancement and audio source separation tasks [109, 84].
III.3.2 BayesWave Comparison
This experiment compares DeepExtractor with BayesWave, the current state-of-the-art method for glitch mitigation in GW physics. Following the identification of the optimal DeepExtractor configuration in Section III.3.1, we conduct a controlled evaluation of both methods using simulated glitch injections.
In this experiment, we inject 30 samples from each of the 12 glitch classes mentioned in Table 1 (360 samples in total) into background noise at SNRs $\in[15,100]$, then process the samples using both DeepExtractor and BayesWave for a direct, 1:1 comparison. We increase the lower SNR bound from $7.5$ (as used in Section III.3.1) to $15$ to prevent potential convergence issues with BayesWave. The performance of each method is assessed by evaluating the mismatch (Eq. 7) between the injected and reconstructed samples, following the same approach as in Section III.3.1.
III.3.3 Gravity Spy Glitches
The final experiment of this study evaluates the performance of our model on Gravity Spy glitches. We conduct a qualitative assessment of Gravity Spy Q-scans before and after glitch subtraction to visually inspect the effectiveness of our model, while also inspecting the resulting glitch and residual noise. Since our fine-tuned DeepExtractor model was trained on 222\,2s samples extracted from 141414\,14s whitened segments, we need to use 141414\,14s of real data surrounding the glitch during inference to ensure the whitening filter remains consistent (see Section III.2.2). To avoid inadvertently suppressing the glitch, we use the PSD from an adjacent 141414\,14s segment that does not include the glitch. This ensures that the glitch is not included in the PSD calculation, while still allowing us to whiten the data effectively. Although the adjacent 141414\,14s segment is not vetoed for other glitches (which could result in imperfect whitening), this approach helps prevent the desired glitch from being suppressed during the whitening process.
For a limited quantitative assessment of the model’s performance on real glitch data, we use the Shapiro-Wilk test [110]. This statistical test compares the p-values of real glitch events before and after mitigation with DeepExtractor. Samples with p-values below a confidence interval (5%) indicate strong evidence for a non-Gaussian component, which is indicative of a glitch. We present the Shapiro-Wilk results in Appendix VII.4, as this test is not a perfect method for characterizing glitches (see Section IV.3 for more details).
IV Results
IV.1 Simulated Experiments
[…]
chirp $2.0^{+2.1}_{-0.8}$ $26.1^{+20.8}_{-13.9}$ $2.3^{+2.4}_{-0.9}$ $41.2^{+20.1}_{-9.3}$ $1.9^{+1.7}_{-0.6}$ $1.1^{+0.9}_{-0.4}$ $\mathbf{0.8}^{+0.7}_{-0.3}$
sine $1.7^{+2.8}_{-0.6}$ $23.5^{+27.6}_{-11.5}$ $1.9^{+3.7}_{-0.7}$ $41.1^{+20.2}_{-9.7}$ $1.7^{+1.8}_{-0.6}$ $0.8^{+1.0}_{-0.2}$ $\mathbf{0.5}^{+0.7}_{-0.1}$
sine_gaussian $1.4^{+1.8}_{-0.6}$ $8.9^{+13.5}_{-3.3}$ $1.6^{+2.0}_{-0.7}$ $40.2^{+16.3}_{-11.3}$ $1.7^{+1.8}_{-0.8}$ $0.6^{+0.7}_{-0.2}$ $\mathbf{0.4}^{+0.6}_{-0.1}$
gaussian_pulse $0.4^{+0.4}_{-0.2}$ $3.8^{+3.8}_{-1.7}$ $0.6^{+0.7}_{-0.3}$ $37.7^{+13.9}_{-11.5}$ $1.8^{+2.1}_{-0.8}$ $0.6^{+0.7}_{-0.2}$ $\mathbf{0.4}^{+0.4}_{-0.2}$
ringdown $0.6^{+0.8}_{-0.2}$ $5.2^{+4.3}_{-2.1}$ $0.9^{+1.0}_{-0.4}$ $39.3^{+15.0}_{-9.5}$ $2.2^{+2.4}_{-1.0}$ $0.9^{+0.8}_{-0.4}$ $\mathbf{0.7}^{+0.6}_{-0.3}$
gengli_H1 $1.0^{+1.0}_{-0.3}$ $3.8^{+3.3}_{-1.4}$ $1.2^{+1.1}_{-0.4}$ $36.6^{+13.0}_{-9.0}$ $2.7^{+2.0}_{-1.1}$ $1.2^{+0.8}_{-0.4}$ $\mathbf{0.9}^{+0.7}_{-0.3}$
gengli_L1 $1.2^{+0.7}_{-0.4}$ $4.0^{+2.8}_{-1.1}$ $1.3^{+0.9}_{-0.4}$ $36.1^{+10.9}_{-7.6}$ $2.8^{+2.1}_{-1.1}$ $1.2^{+0.9}_{-0.3}$ $\mathbf{1.1}^{+0.7}_{-0.4}$
cdvgan_blip $1.2^{+1.0}_{-0.4}$ $4.5^{+3.9}_{-1.7}$ $1.4^{+1.1}_{-0.5}$ $36.2^{+15.6}_{-9.1}$ $2.9^{+2.6}_{-1.0}$ $1.2^{+0.9}_{-0.4}$ $\mathbf{0.9}^{+0.8}_{-0.2}$
cdvgan_tomte $1.0^{+0.7}_{-0.3}$ $4.5^{+3.4}_{-1.7}$ $1.2^{+0.8}_{-0.4}$ $38.0^{+12.4}_{-10.7}$ $2.6^{+2.5}_{-1.0}$ $1.0^{+0.7}_{-0.3}$ $\mathbf{0.9}^{+0.6}_{-0.3}$
cdvgan_bbh $1.0^{+1.1}_{-0.3}$ $4.6^{+5.2}_{-1.5}$ $1.1^{+1.2}_{-0.3}$ $34.6^{+17.7}_{-6.7}$ $2.7^{+2.4}_{-0.9}$ $1.2^{+1.0}_{-0.5}$ $\mathbf{0.8}^{+0.9}_{-0.2}$
cdvgan_simplex $1.8^{+1.4}_{-0.6}$ $5.2^{+5.4}_{-1.6}$ $2.0^{+1.6}_{-0.7}$ $36.9^{+16.1}_{-7.6}$ $3.5^{+3.5}_{-1.1}$ $1.8^{+1.4}_{-0.5}$ $\mathbf{1.6}^{+1.2}_{-0.5}$
cdvgan_uniform $1.6^{+1.1}_{-0.5}$ $5.3^{+3.9}_{-1.9}$ $1.7^{+1.3}_{-0.5}$ $38.1^{+13.3}_{-10.0}$ $3.1^{+3.3}_{-1.1}$ $1.6^{+1.0}_{-0.5}$ $\mathbf{1.4}^{+0.9}_{-0.5}$
Total $1.2^{+1.1}_{-0.4}$ $5.9^{+6.9}_{-2.4}$ $1.4^{+1.3}_{-0.5}$ $37.9^{+15.2}_{-9.2}$ $2.5^{+2.3}_{-0.9}$ $1.1^{+0.9}_{-0.4}$ $\mathbf{0.9}^{+0.7}_{-0.3}$
Table 1: Median mismatch (%) between injected and extracted glitches over 512 samples from each class. The bounds represent the range of the $1\sigma$ credible interval. Bold text highlights the best model for each signal type.
Figure 7: DeepExtractor reconstructions (orange) of samples comprising multiple injections (black). The upper panels display the full $h(t)$ signals (gray), while the lower panels zoom in on a 0.50.50.5\,0.5s segment (outlined with a red box) to highlight interesting features of the reconstructed injections after background removal.
Table 1 presents the median mismatch scores for each deep learning model across our test dataset, which consists of 512 samples from each class (totaling 6,144 test samples). The bounds represent the $1\sigma$ credible interval. All models in the table adhere to the framework outlined in FIG. 2, where the background noise, $n(t)$, is first predicted and then subtracted from the input to reconstruct the glitch. The mismatch distributions are generally skewed towards low values for most models. However, a small subset of the samples includes very low SNR injections, located at the lower end of the SNR range (7.5). These low-SNR injections subtly affect the background noise, resulting in significantly higher mismatch results for these specific cases across all models. This trend is evident in the upper bounds shown in Table 1, which are consistently larger than the corresponding lower bounds.
[…]
Although slightly trailing DeepExtractor in performance, the time-domain variant, UNET1D, still delivers impressive results, achieving a median mismatch of 1.2%. Similarly to DeepExtractor, mapping directly to $g(t)$ with UNET1D does not improve the performance of mapping to $n(t)$, where a median mismatch of $1.3^{+1.2}_{-0.5}\%$ is obtained across the test dataset. This highlights the utility of our framework across time and spectrogram domains. Furthermore, using both the dual-channel output and difference layer configuration make no notable improvement, with both yielding a median mismatch of 1.3%.
In contrast to the Autoencoder1D model, which produces mismatch results similar to UNET1D, the Autoencoder2D model exhibits notably large mismatch values between the reconstructed and injected glitches. This suggests that feature compression within the Autoencoder layers—particularly in the spectrogram domain—may be a critical issue. To explore this further, we removed one of the Autoencoder2D layers to assess whether reduced compression could lead to better results in the spectrogram domain, but this adjustment did not improve the results beyond those presented in Table 1. This finding highlights the importance of skip connections in U-Nets, which play a vital role in preserving key features and enhancing reconstruction accuracy.
Finally, applying DeepExtractor to 512 pure background samples without injections (i.e., where it is expected to reproduce the input) yields a maximum mismatch of $0.26\%$ between the target and the reconstructed background noise. This demonstrates the model’s ability to accurately replicate the background noise in the absence of signals or glitches, ensuring that no extraneous signals or glitches are reconstructed when they are not present in the data.
FIG. 7 shows reconstructions of multi-injection samples, similar to the training data, demonstrating DeepExtractor’s versatility. It consistently recovers diverse glitches, even when heavily obscured by noise. Across 512 samples, DeepExtractor achieves a median mismatch of $0.7^{+0.2}_{-0.2}$. The model excels with multi-injection samples by effectively matching non-Gaussian components, which are more frequent per sample and better aligned with the training data than the single-class samples in Table 1555BayesWave was not applied to these samples, as it requires a single trigger time, while these samples have multiple trigger times. We instead maintain a controlled comparison with single-class samples, where results are easier to interpret..
IV.2 BayesWave Comparison
FIG. 8 presents a comparison of mismatch performance between DeepExtractor and BayesWave on a reduced dataset, consisting of 30 samples per class. For reference, examples of reconstructions from both BayesWave and DeepExtractor are shown in FIG. 9. DeepExtractor consistently outperforms BayesWave, even on previously unseen test classes. Specifically, DeepExtractor achieves a median mismatch of $0.3^{+0.2}_{-0.1}\%$ across the entire test dataset, showing a significant improvement over BayesWave, which has a median mismatch of $1.9^{+2.2}_{-1.1}\%$ (the bounds represent the $1\sigma$ confidence interval). The performance improvement of DeepExtractor compared to Section IV.1 can be attributed to the lower SNR limit being raised to 15, making the reconstruction task easier.
The most notable improvements are observed in the chirp, sine, sine-gaussian, and ringdown classes, where BayesWave reaches maximum mismatches as high as $81.8\%$, $66.8\%$, $72.0\%$, and $83.9\%$, respectively. In contrast, DeepExtractor shows much lower maximum mismatches of $4.0\%$, $3.0\%$, $1.9\%$, and $16.5\%$ for these classes.
The discrepancies may be due to BayesWave’s difficulty in converging due to high-frequency features in the these samples. Furthermore, DeepExtractor demonstrates significant improvements when applied to realistic glitch samples from glitch generators, particularly hybrid samples from cDVGAN, which feature diverse glitch morphologies.
Finally, DeepExtractor also offers a substantial advantage in terms of speed, reconstructing a single glitch sample in an average of $0.14$ seconds on a CPU (or $0.05$ per glitch seconds when processing a batch of 256 samples). This represents a speedup of over $10,000$ times compared to BayesWave, which relies on a traditional Bayesian framework.
Figure 8: Box plots showing mismatch between injected and reconstructed samples per class. Lower boxplots for DeepExtractor indicate superior glitch reconstruction performance.
Figure 9: Examples from our comparison between BayesWave and DeepExtractor. Injected SNRs and mismatches yielded from both approaches are shown above each plot.
IV.3 Gravity Spy Glitches
(a) Input Glitch
(b) Without transfer learning
(c) With transfer learning
(d) Time-series outputs
Figure 10: An example of subtracting a blip glitch with and without transfer learning DeepExtractor on real detector noise. FIG. 10(b) shows that, without transfer learning on real detector noise, the line features at 512512512\,512Hz are also subtracted, highlighted within the red lines. This is because the model was trained on a purely flat, simulated PSD, causing the realistic line features to also be reconstructed as part of the excess power.
This experiment applies transfer learning to fine-tune DeepExtractor on real LIGO O3 data from both the Hanford and Livingston detectors. The primary objective is to improve the model’s ability to reconstruct real glitch events. Given the complex, unmodeled nature of glitches, we take a qualitative approach by visualizing DeepExtractor’s reconstruction performance through Q-scan images and time-series plots.
FIG. 10 illustrates the reconstruction of a blip glitch from LIGO Hanford during the O3a run, before and after transfer learning with DeepExtractor. This example highlights the significance of further training with real detector data, as it includes features such as high-frequency lines that are common in actual LIGO data. Incorporating these realistic features is crucial for the model to accurately differentiate between detector noise and genuine glitch features, improving its reconstruction capabilities.
FIG. 10(d) displays the time-domain output of both the pre-trained (blue) and fine-tuned (orange) models. Without transfer learning, the model erroneously reconstructs the line features, obscuring the true glitch signal. However, after transfer learning, the model has learned to properly identify and reconstruct these features, resulting in a more accurate glitch reconstruction.
Additional examples of DeepExtractor’s fine-tuned performance on Gravity Spy glitches are provided in Appendix VII.5. The Q-scans show that, for most glitch classes, noise features are effectively preserved after glitch removal. However, for louder glitches, such as those from the ‘Extremely Loud’ and ‘Koi Fish’ classes, where the glitch dominates the signal, the removal process leads to a“zeroing out” effect on the data, as seen in the Q-scans post-mitigation. However, in most cases, the time-series plots demonstrate that the reconstructions align closely with the actual data. Overall, DeepExtractor exhibits flexibility across glitch classes.
Shapiro-Wilk test results, presented in Appendix VII.4, show that most glitch classes have average p-values above the 5% significance threshold, suggesting limitations in the Shapiro-Wilk test for characterizing glitches. However, higher p-values after glitch mitigation indicate that the residual data better approximates a Gaussian distribution, signaling that DeepExtractor has improved the data quality.
V Conclusions and Future Work
This paper introduced DeepExtractor, a novel deep learning framework designed to reconstruct signals and glitches in gravitational wave detector data. Leveraging a PSD-informed U-Net architecture, DeepExtractor processes magnitude and phase spectrograms derived from the short-time Fourier transform (STFT). This spectrogram-based approach enables transformations between the frequency and time domains, essential for tasks such as glitch mitigation and signal reconstruction.
The key innovation lies in the training strategy. Rather than directly mapping input data to clean waveforms, the model is trained to isolate the noise component from input data comprising background noise and injected signals or glitches. By subtracting the predicted noise, DeepExtractor reconstructs the signal or glitch. This approach outperforms direct mapping in generalization and underscores the importance of accurately modeling the detector’s power spectral density (PSD). The framework is highly flexible, capable of reconstructing any power excess above the inherent detector noise. This flexibility stems from training on a diverse dataset of linear combinations of waveforms from five distinct analytical classes, enabling effective interpolation to reconstruct unseen waveforms.
Three experiments validated DeepExtractor’s efficacy. The first experiment involves simulated Gaussian noise with injected glitches, measuring time-domain match performance across several deep learning models. The results demonstrate that our U-Net architecture significantly outperforms autoencoders without skip connections, even when the overall architectures are otherwise identical. Additionally, U-Nets operating on magnitude and phase spectrograms are shown to deliver superior performance compared to a 1D U-Net applied directly to the corresponding time-series data.
In the second experiment, we compared DeepExtractor to the state-of-the-art BayesWave algorithm in a simulated environment. Our results reveal that DeepExtractor not only delivers higher accuracy and greater stability than BayesWave but also demonstrates a key advantage of deep learning over traditional Bayesian approaches: computational efficiency. While BayesWave relies on computationally intensive Bayesian inference, DeepExtractor achieves a remarkable speedup of over 10,000 times on a CPU, demonstrating its potential for real-time gravitational wave data analysis.
Finally, we evaluated DeepExtractor’s capability to reconstruct real glitch events from LIGO’s third observing run. This represents the first comprehensive application of deep learning to glitch reconstruction. By utilizing Q-scan and time-domain visualizations, we ensured effective glitch removal while preserving the underlying noise characteristics. The study highlighted the importance of integrating PSD knowledge through transfer learning, allowing the model to adapt effectively to the complexities of real detector data.
For signal reconstruction, a key next step is to compare DeepExtractor’s performance against established search and parameter estimation pipelines. Currently, the model does not handle overlapping signals and glitches, but this limitation could be addressed by extending the training framework. Instead of isolating only the noise component, the model could be trained to separate noise and signals from data containing noise, signals, and glitches. Simulated signal components could be generated using waveform models commonly employed in parameter estimation. Incorporating data from multiple detector channels could further enhance performance, particularly in scenarios with overlapping signals and glitches.
Recent research leveraging the null stream in the Einstein Telescope’s triangular design has shown effective glitch mitigation in the presence of overlapping signals, marking a significant advancement for the third-generation detector era [111]. Building on this progress, DeepExtractor could further enhance the null stream’s performance by precisely isolating uncorrelated glitches from signals. Integrating DeepExtractor into this framework would be a straightforward process and could strongly reinforce the scientific case for the Einstein Telescope’s triangular design.
VI Acknowledgements
This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by the National Science Foundation Grants No. PHY-0757058 and No. PHY-0823459.
VII Appendixes
VII.1 Simulated training glitches
Table 2 outlines the analytical models employed to generate the proxy glitch classes for training DeepExtractor, along with the parameters used in their functions.
Function Description Key Parameters
chirp Generates a linear chirp signal with frequency varying from f0 to f1 over the duration. f0_min, f0_max, f1_min, f1_max, duration, sample_rate
sine Generates a simple sine wave with a frequency randomly chosen between freq_min and freq_max. freq_min, freq_max, duration, sample_rate
sine-gaussian Generates a sine wave modulated by a Gaussian envelope. Frequency and tau are randomly chosen. freq_min, freq_max, duration, sample_rate
Gaussian pulse Generates a Gaussian pulse with randomized frequency and bandwidth characteristics. fc_min, fc_max, bw_min, bw_max, bwr_min, bwr_max, tpr_min, tpr_max, duration, sample_rate
ringdown Generates a ringdown signal based on a damped sinusoidal model with a random frequency f0, decay time tau, and quality factor Q. There is also a 50% probability to flip the signal horizontally. duration, sample_rate, n_signals
Table 2: Summary of analytical glitch models used to simulate glitches for training DeepExtractor.
VII.2 STFT Parameters
Table 3 details the parameters and their respective values used for STFT computation, enabling transformations to and from the time-frequency (spectrogram) domain.
Parameter Description Value
n_fft The size of the Fast Fourier Transform (FFT), typically a power of 2. This determines the resolution of the frequency analysis. 512
window_length The length of the window function used in the FFT. This determines how much of the signal is considered at once. 64
hop_length The step size between consecutive windows when performing the FFT. It defines how much overlap there is between consecutive windows. 32
Hann window A window function often used in FFT analysis due to its smooth tapering, reducing spectral leakage. It satisfies the COLA condition when paired with a hop length of half the window length. Used
[…]
Table 3: Overview of parameters used in the short-time Fourier transform (STFT) for transforming between the time and spectrogram domains in DeepExtractor.
VII.3 Reconstruction of unseen test classes
FIG. 11 shows examples of full 222\,2s extracted signals and zoomed-in views (middle 0.50.50.5\,0.5s) for classes that were unseen during training.
(a) Gengli H1
(b) cDVGAN Tomte
(c) cDVGAN BBH
(d) cDVGAN Simplex
Figure 11:
VII.4 Results of Shapiro Tests on Gravity Spy data
Table 4 shows the mean and standard deviation of Shapiro-Wilk test p-values before and after mitigation via DeepExtractor over 100 random samples per class. Higher p-values after analyzing with DeepExtractor indicate that the output looks more Gaussian.
Category Before After
Extremely_Loud 0.00 $\pm$ 0.00 0.39 $\pm$ 0.33
Scattered_Light 0.49 $\pm$ 0.30 0.52 $\pm$ 0.29
Scratchy 0.36 $\pm$ 0.31 0.46 $\pm$ 0.27
Power_Line 0.43 $\pm$ 0.33 0.50 $\pm$ 0.32
Koi_Fish 0.00 $\pm$ 0.00 0.46 $\pm$ 0.30
Whistle 0.26 $\pm$ 0.34 0.50 $\pm$ 0.31
Low_Frequency_Burst 0.46 $\pm$ 0.30 0.47 $\pm$ 0.29
Tomte 0.16 $\pm$ 0.23 0.45 $\pm$ 0.28
Repeating_Blips 0.04 $\pm$ 0.17 0.44 $\pm$ 0.32
Fast_Scattering 0.49 $\pm$ 0.31 0.50 $\pm$ 0.30
Blip 0.03 $\pm$ 0.12 0.49 $\pm$ 0.28
Blip_Low_Frequency 0.37 $\pm$ 0.31 0.42 $\pm$ 0.31
Violin_Mode 0.13 $\pm$ 0.23 0.39 $\pm$ 0.27
Helix 0.02 $\pm$ 0.10 0.41 $\pm$ 0.31
Low_Frequency_Lines 0.47 $\pm$ 0.31 0.47 $\pm$ 0.31
Wandering_Line 0.17 $\pm$ 0.24 0.35 $\pm$ 0.33
No_Glitch 0.38 $\pm$ 0.30 0.41 $\pm$ 0.30
Air_Compressor 0.52 $\pm$ 0.28 0.53 $\pm$ 0.28
Chirp 0.24 $\pm$ 0.31 0.40 $\pm$ 0.28
Light_Modulation 0.00 $\pm$ 0.02 0.54 $\pm$ 0.28
Paired_Doves 0.29 $\pm$ 0.32 0.47 $\pm$ 0.30
1400Ripples 0.37 $\pm$ 0.35 0.49 $\pm$ 0.31
1080Lines 0.52 $\pm$ 0.32 0.57 $\pm$ 0.31
Table 4: Mean Shapiro-Wilk test p-values before and after glitch mitigation via DeepExtractor.
VII.5 Extracted Gravity Spy Glitches
Figures 12, 13, 14, 15 and 16 show DeepExtractor reconstructions for glitch classes from the Gravity Spy dataset.
(a) Air Compressor (03a, L1)
(b) Blip (03b, H1)
(c) Blip Low Freq. (03b, L1)
Figure 12:
(a) Extremely Loud (03a, H1)
(b) Helix (03b, L1)
(c) Koi Fish (03b, L1)
Figure 13:
(a) Light Modulation (03a, L1)
(b) Low Frequency Burst (03b, H1)
(c) Paired Doves (03b, H1)
Figure 14:
(a) Power Line (03a, L1)
(b) Repeating Blips (03a, H1)
(c) Scattered Light (03a, H1)
Figure 15:
(a) Scratchy (03b, L1)
(b) Wandering Line (03b, H1)
(c) Whistle (03a, H1)
Figure 16: